OpenStudy (goformit100):

A and B are two independent events. The probability that both A and B occur is 1/6 and the probability that neither of them occurs is 1/3. Find the probability of the occurrence of A.

OpenStudy (goformit100):

@genius12

OpenStudy (anonymous):

P(A int B) = P(A) * P(B)

OpenStudy (goformit100):

How ?

OpenStudy (anonymous):

since the events are independent..

OpenStudy (anonymous):



OpenStudy (goformit100):

oh wao. @satellite73 you gave me the clear conception. and hey that's calles a True Experienced Guru :) Thankyou Sir.

OpenStudy (goformit100):

*called

OpenStudy (goformit100):

your way of explaning me using venn-diagram gave me help

OpenStudy (anonymous):

well we are not done just by the picture you have to use independence as well

OpenStudy (goformit100):

ya I know

OpenStudy (anonymous):

\[P(A)+P(B)-P(A\cap B)=P(A\cup B)=\frac{2}{3}\]

OpenStudy (anonymous):

so \(P(A)+P(B)=\frac{5}{6}\)

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

i think there is no other way but to solve \[P(A)P(B)=\frac{1}{6}, P(A)+P(B)=\frac{5}{6}\]

OpenStudy (goformit100):